if a plane is travelling 3100 km at 790km/h and then encounters a tailwind that boosts its speed tp 990km/h for the next 2800km. what was the total time for the trip?what was teh average speed of the plane for this trip?
Answer by ptaylor(1894)

The plane travels 3100 km (d1) at 790 km/h (r1) and 2800 km (d2) at 990km/hr (r2)
Let t1=time it takes the plane to travel 3100km @790km/hr
and t2=time it takes the plane to travel 2800km @990km/hr
We'll use the formula:
distance=(rate)(time) or d=rt
t1=d1/r1 =(3100 km)/(790km/hr)=3.92 hrs and
t2=d2/r2 =(2800 km)/(990km/hr)=2.83 hrs
Total time for the trip =t1+t2=2.83 hrs+3.92 hrs =6.75 hrs
average rate = (total distance)/(total time)
total distance =5900 km
total time=6.75 hrs

average rate=5900/6.75=874.1 km/hr
ck
d=rt
5900 km=(874.1km/hr)(6.75 hr)
5900 km =5900 km
Hope this helps. Happy holidays---ptaylor